Archive work on 22 October 2009 UT

The LBC data flow was changed (Riccardo & Dave) so that the data go directly to /newdata which is mounted on the observer workstation, to /Repository/UTdate, also mounted on the observer workstations. This change has necessitated changes to:
  • the partner accounts' .cshrc files:
setenv LBCFPIA_DATADIR "/newdir"
  • the scripts which automatically display data and list files: quickfile.py, Blue_Science.py and Red_Science.py. I simply changed Repository to newdata.

Plan for the 2 tech nights following the LBC science time

1. Calibrate PSF aberrations and compare FPIA with EF Collimate, manually add Z11, take extra/intra focal pair with RB_rVexin4.ob, manually remove Z11 and collimate again, manually add Z11, run RB_rVexin4.ob, manually remove, etc...
Z11 = -2000, +2000, -1500, +1500, -1000, +1000, -500, +500, 0 <br>
Z22 = -300, +300, -200, +200, -100, +100, 0. <br>
Z7 = -2000, +2000, -1500, +1500, -1000, +1000, -500, +500, 0. <br>
Z5 = -2000, +2000, -1500, +1500, -1000, +1000, -500, +500, 0. <br>
2. Calibrate PSF with EF, using larger pupils of 100 pix diameter as suggetsed for EF Take extra/intra focal pairs with larger focus offset than 0.8mm --- aim for 100 pixels diameter so +/- 1.3mm should give that (/home/LBTO/Calib_OBs/FOCUS/RB_rV100exin4.ob). Do for Z4, but if there is more time, the same could be done with Z5, Z11 and Z22 (+/-200).
collimate
give Z4 = -1000,
run RB_rV100exin4.ob
remove Z4 and collimate again
give Z4= +1000
run RB_rV100exin4.ob
remove Z4 and collimate again

3. Assess the effect of seeing on measurements of aberrations, particularly Z22 Take extra/intra focal pairs at a range of seeing values: Collimate, extra/intra focal (RB_rVexin4.ob)

4. Check focus corrections from tech chips --- focus/collim, observe an OB, focus/collim again. How many iterations does it take?


1. Red master/ Blue slave issue trak #1989. Insure that the instructions for switching - running Blue as master when Red is not used, work. Also insure that Red remains master when Blue is not used. Address Norm's email. 2. Correct co-pointing --- determine what global offsets blue needs to copoint at high elevation. 3. Filter focus offsets - R-BESSEL, i-SLOAN, I-BESSEL used tonight (20091020) and esp R-BESSEL seemed off at the beginning of the night - or was it the focus/collim? 4. Noise and high background in LBC-Red.

Notes: GUI should grey out 2x2 binning for Blue & address Red/Blue master/slave hardwiring.

-- OlgaKuhn - 22 Oct 2009
Topic revision: r2 - 23 Oct 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback